# LAGUNA: LAnguage Guided UNsupervised Adaptation with structured spaces

### v3

@@ 3.3 Stage 2: Training of the Language Supervisor
Similar to [27], we train a language model to map image captions to semantic labels, with the aim to provide pseudo-labels for target samples, which are associated with captions but not class labels. Differently from [27], we use the language supervisor both to provide pseudo-labels for target samples, useful to train the classifier, and textual representations semantically structured as the domain-agnostic anchors $\mathcal{A}$, useful to encourage alignment to the reference structure. Hence, rather than training a regular classifier, we directly supervise a language model $\mathcal{G}$ to provide representations that are 1) geometrically aligned to $\mathcal{A}$ and 2) suitable for predicting class labels. Specifically, given a pair of source caption and corresponding label $(l_{i}^{s},y_{i}^{s})$, $\mathcal{G}$ processes $l_{i}^{s}$ to produce a vector representation $z^{s}_{i}$:
$z_{i}^{s}=\mathcal{G}(l_{i}^{s}).$ (1)
Next, we encourage the vector representation $z_{i}^{s}$ to be geometrically aligned to $\mathcal{A}[y_{i}^{s}]$ corresponding to ground truth action $y_{i}^{s}$. To do so, we compute $\mathbf{r}^{z_{i}^{s}}=rel(z_{i}^{s},\mathcal{A})$, the relative encoding of $z_{i}^{s}$, and supervise it to be similar to $\mathbf{r}^{y_{i}^{s}}=rel(\mathcal{A}[y_{i}^{s}],\mathcal{A})$, the relative encoding of the anchor $\mathcal{A}[y_{i}^{s}]$, with the following structure-preserving loss:
@@
This loss encourages the encodings $z_{i}^{s}$ of text descriptions $l_{i}^{s}$ associated with labels $y_{i}^{s}$ to preserve the same geometrical associations as their corresponding anchor, encouraging the latent space learned by $\mathcal{G}$ to mirror the structure defined by $\mathcal{A}$. To further favor alignment to $\mathcal{A}$, rather than employing a classification head, we predict class probabilities directly by Softmax-normalizing relative encodings:
@@ 3.4 Stage 3: Cross-Domain Visual Classifier
Stage 3 trains a cross-domain visual classifier aligning representations extracted through a visual encoder $\mathcal{V}$ to the structure imposed by the domain-agnostic language anchors $\mathcal{A}$. We allow each domain to develop its own latent space, but we encourage both spaces to be aligned to two sets of learnable anchors $\mathcal{A}_{t}\in\mathbb{R}^{N_{c}\times D_{v}}$ and $\mathcal{A}_{s}\in\mathbb{R}^{N_{c}\times D_{v}}$ specific to the target and source domain respectively. We further ensure that the structures of $\mathcal{A}_{t}$ and $\mathcal{A}_{s}$ are aligned to the structure of $\mathcal{A}$ through the supervision provided by ground truth labels $y_{i}^{s}$ in the source domain and the text supervision provided by $\mathcal{G}$ in the target domain in the form of the predicted vector $z_{i}^{t}$ and the predicted pseudo-label $\overline{y}_{i}^{t}$. The source and target datasets are merged ($\hat{\mathcal{X}}=\mathcal{X}_{t}+\mathcal{X}_{s}$) and used for training with a total of $M=N_{s}+N_{t}$ samples. Source samples comprise an image, caption, and label triplet $(x_{i}^{s},l_{i}^{s},y_{i}^{s})$, whereas target samples $(x_{i}^{t},l_{i}^{t},\overline{y}_{i}^{t})$ replace label $y_{i}^{s}$ with the pseudo-label $\overline{y}_{i}^{t}$. Target and source images are processed by $\mathcal{V}$ to obtain latent representations $g_{i}^{t}$ and $g_{i}^{s}$, while target captions $l_{i}^{t}$ are processed by $\mathcal{G}$ to obtain latent representations $z_{i}^{t}$ as follows:
@@
These representations and the anchors are used to achieve two main training objectives: 1) structure learning and 2) classification training.
@@ 3.4.1 Structure Learning
To ensure the learned visual representations adhere to the structure defined by $\mathcal{A}$, LAGUNA employs a supervised learning approach based on relative encodings. For target domain samples (without ground truth labels), we first compute the relative encoding $\mathbf{r}^{z_{i}^{t}}=rel(z_{i}^{t},\mathcal{A})$. For source domain samples, we compute the relative encoding of the anchor associated to the ground truth class $\mathcal{A}[y_{i}^{s}]$: $\mathbf{r}^{y_{i}^{s}}=rel(\mathcal{A}[y_{i}^{s}],\mathcal{A})$. These encodings represent the relative positions of textual representations with respect to the reference space $\mathcal{A}$ and are not trainable as $\mathcal{G}$ is fixed. On the visual side, we compute the relative encodings of $g_{i}^{t}$ and $g_{i}^{s}$ with respect to the learnable anchors $\mathcal{A}_{t}$ and $\mathcal{A}_{s}$, i.e., $\mathbf{r}^{g_{i}^{t}}=rel(g_{i}^{t},\mathcal{A}_{t})$ and $\mathbf{r}^{g_{i}^{s}}=rel(g_{i}^{s},\mathcal{A}_{s})$. We encourage the visual relative encodings ($\mathbf{r}^{g_{i}^{t}},\mathbf{r}^{g_{i}^{s}}$) to match the text relative encodings ($\mathbf{r}^{z_{i}^{t}},\mathbf{r}^{y_{i}^{s}}$) with an L1 loss:
@@
This loss encourages visual representations to maintain relationships with their respective domain-specific visual anchors that mirror the relationships expressed by the domain-invariant anchors $\mathcal{A}$. This process supervises both the learnable anchors $\mathcal{A}_{s},\mathcal{A}_{t}$ and the visual encoder $\mathcal{V}$. However, since this loss alone might lead to the collapse of anchor representations into a small region, hindering effective classification, we introduce a volume (or spread) regularization loss. This loss treats each set of anchors as a multidimensional parallelotope whose volume in the latent space can be measured by the determinant of its Gram matrix. Given the Gram matrices for the three sets of anchors:
@@
we encourage the volume of each domain-specific parallelotope to be approximately equal to the volume of the reference anchors $\mathcal{A}$ with the following regularization loss:
$\mathcal{L}_{Reg}$ $=$ $|logDet(\gamma_{t})-logDet(\gamma)|+$ (8)
$|logDet(\gamma_{s})-logDet(\gamma)|,$
where we use log-determinants for numerical stability. This regularization loss ensures taht visual anchors occupy a volume similar to the one of $\mathcal{A}$, preventing collapse and promoting better class separation in the representation space.
@@ 3.4.2 Classification Training
We perform classification training concurrently with structure learning. To ensure that source and target visual representations are grounded in the structure imposed by $\mathcal{A}$, while still being free to capture domain-specific nuances, we propose a novel cross-domain attention layer that aims to ground domain-specific visual representations $g_{i}^{t}$ and $g_{i}^{s}$ to the structure of source anchors $\mathcal{A}_{s}$. We include two Cross-Domain attention layers sharing the same weights for the target and source branches of LAGUNA (See Fig. 2). In the target branch, we use $g_{i}^{t}$ as queries, target anchors $\mathcal{A}_{t}$ as keys, and source anchors $\mathcal{A}_{s}$ as values. The output is summed to $g_{i}^{t}$ with a residual connection to include both domain-specific and cross-domain information in the final representation:
@@
Crucially, the cross-domain attention layer constructs an attention map by leveraging the domain-specific relations between the encoded representations and the anchors (i.e., $QK^{T}$). These relations are expected to be similar across domains due to the influence of the structural loss $\mathcal{L}_{S}$. This process results in a unified representation since the output of the attention layer always depends on values coming from the source anchors $\mathcal{A}_{s}$. The residual connection included in the attention layer is introduced to account for domain-dependent class characteristics, which results in the feature vectors $f_{i}^{s}$ and $f_{i}^{t}$. Finally, an MLP classification head processes $f_{i}^{t}$ and $f_{i}^{s}$ to output predictions $\hat{y}_{i}^{t}$ and $\hat{y}_{i}^{s}$. The MLP has shared weights across domains and is optimized with a cross-entropy loss using the ground truth label $y_{i}^{s}$ for source examples and the pseudo-label $\overline{y}_{i}^{t}$ predicted by $\mathcal{G}$ in the case of target samples. The overall training objective is:
$\small\mathcal{L}=\lambda_{1}\mathcal{L}_{CE}(\hat{y_{i}},y_{i}^{*})+\lambda_{2}\mathcal{L}_{S}(\mathbf{r}^{*},\mathbf{r}^{g_{i}})+\lambda_{3}\mathcal{L}_{Reg}(\gamma,\gamma_{*}),$ (11)
@@ 4 Experiments
MDD [77] 52.80 41.20 47.80 52.50 42.10 40.70 54.20 54.30 43.10 51.20 43.70 41.70 47.11
SCDA [34] 54.00 42.50 51.90 55.00 44.10 39.30 53.20 55.60 44.70 56.20 44.10 42.00 48.55
SSRT-B [64] 69.90 58.90 66.00 75.80 59.80 60.80 73.20 70.60 62.20 71.40 61.70 55.20 65.41
MemSAC [25] 63.49 42.14 60.32 72.33 54.92 46.14 73.46 68.04 52.75 74.42 57.79 43.57 59.11
CDTrans [75] 66.20 52.90 61.50 72.60 58.10 57.10 72.50 69.00 59.00 72.10 62.90 53.90 63.16
PMTrans [79] 74.10 61.10 70.00 79.30 63.70 62.70 77.50 73.80 62.60 79.80 69.70 61.20 69.63
@@
TextMatch [27] 71.36 64.30 65.32 81.25 65.65 64.85 81.09 72.65 63.94 81.08 70.84 64.17 70.64
nGramMatch [27] 68.92 59.82 63.15 76.35 61.72 62.87 76.35 69.28 62.51 76.04 68.52 60.52 67.17
LaGTran [27] 77.30 68.25 67.35 81.31 67.03 66.81 80.78 75.62 68.08 79.23 73.80 63.44 72.41
@@
ViT-B CLIP* [56] 72.39 60.90 66.81 81.37 60.90 66.81 81.37 72.39 66.81 81.37 72.39 69.90 70.38
DAPrompt* [16] 73.90 65.90 70.40 84.90 65.80 70.20 84.60 73.50 69.90 84.90 73.80 65.80 73.80
PADCLIP [31] 76.40 67.50 72.70 84.20 68.10 71.10 83.60 76.30 71.70 83.50 75.40 67.20 74.81
@@
Table 1: Results on DomainNet [52] in $4$ domains: Real, Clipart, Sketch, and Painting, for 12 adaptation scenarios using the Swin-B [39] and ViT-B [12] backbones. The best results are in bold, second bests are underlined.‘*‘ method uses domain-aware prompts, $\dagger$ our run.
4.1 Experimental Setup
Datasets. We evaluate LAGUNA on four comprehensive UDA benchmarks: DomainNet [52], GeoPlaces [26] GeoImnet [26], and Ego2Exo [27, 20]. DomainNet consists of $400K$ images across $345$ classes and is used to evaluate performance in $12$ domain transfer settings [70, 27]. GeoImnet and GeoPlaces are subsets of the larger GeoNet dataset [26], contain over 750K images and focus on geographic disparities in image classification (GeoImNet with $600$ classes) and scene recognition (GeoPlaces with $205$ classes). Finally, Ego2Exo is a subset of EgoExo-4D [20] curated for domain adaptation. It involves a total of $9086$ videos and $24$ action categories. DomainNet and GeoNet include per-image captions derived from BLIP-2 [33] and web metadata, respectively, while Ego2Exo provides video descriptions from the original EgoExo-4D dataset.
Model and training details. We use a pre-trained SentenceTransformer [57] for structure definition (Stage 1) and a BERT base model [61] as text supervisor (Stage 2). The latter is trained on the source domain of each scenario for 5 epochs with a batch size of 64, a lr of 1e-4, and the AdamW [43] optimizer. The visual encoders (Stage 3) are tailored to each dataset for fair comparisons with existing approaches. For DomainNet, we adapt the Swin-B and ViT-B backbones [39, 12]. For GeoImnet and GeoPLaces, we utilize the ViT-B backbone [12] following [70, 79]. For Ego2Exo, we employ pre-extracted Omnivore-L features [20] as in [27]. All vision models undergo joint training with the other network components for $10$ epochs with a batch size of $32$, an initial learning rate of $41e-$ 4, and cosine scheduling. We set the loss weights to $\lambda_{1}$=1.0 $\lambda_{2}$=0.1 and $\lambda_{3}$=0.001 to normalize the relative magnitude of the related loss functions.

### v4

@@ 3.3 Stage 2: Training of the Language Supervisor
Similar to (Kalluri et al., 2024), we train a language model ($\mathcal{G}$) from captions to provide pseudo-labels for target samples, which have no class labels. In addition, we also use $\mathcal{G}$ to learn textual representations semantically structured as the domain-agnostic anchors $\mathcal{A}$, useful to encourage alignment to the reference structure. Hence, rather than training a regular classifier, we directly supervise $\mathcal{G}$ to provide representations that are 1) geometrically aligned to $\mathcal{A}$ and 2) suitable for predicting class labels. Specifically, given a pair of source caption-label $(l_{i}^{s},y_{i}^{s})$, $\mathcal{G}$ processes $l_{i}^{s}$ to produce a vector representation $z^{s}_{i}=\mathcal{G}(l_{i}^{s})$. Next, we encourage the vector representation $z_{i}^{s}$ to be geometrically aligned to $\mathcal{A}[y_{i}^{s}]$ corresponding to ground truth action $y_{i}^{s}$. To do so, we compute $\mathbf{r}^{z_{i}^{s}}=rel(z_{i}^{s},\mathcal{A})$, the relative encoding of $z_{i}^{s}$, and supervise it to be similar to $\mathbf{r}^{y_{i}^{s}}=rel(\mathcal{A}[y_{i}^{s}],\mathcal{A})$, the relative encoding of the anchor $\mathcal{A}[y_{i}^{s}]$, with the following structure-preserving loss:
@@
This loss encourages the encodings $z_{i}^{s}$ of text descriptions $l_{i}^{s}$ associated with $y_{i}^{s}$ to preserve the same geometrical associations as $\mathcal{A}[y_{i}^{s}]$, encouraging the latent space learned by $\mathcal{G}$ to mirror the structure defined by $\mathcal{A}$. To further favor alignment to $\mathcal{A}$, rather than employing a classification head, we predict class probabilities directly by Softmax-normalizing relative encodings:
@@ 3.4 Stage 3: Cross-Domain Visual Classifier
Stage 3 trains a cross-domain visual classifier aligning representations extracted through a visual encoder $\mathcal{V}$ to the structure imposed by $\mathcal{A}$. We allow each domain to develop its own latent space, but we encourage both spaces to be aligned to two sets of learnable anchors $\mathcal{A}_{t}\in\mathbb{R}^{N_{c}\times D_{v}}$ and $\mathcal{A}_{s}\in\mathbb{R}^{N_{c}\times D_{v}}$ specific to the target and source domain respectively. We further ensure that the structures of $\mathcal{A}_{t}$ and $\mathcal{A}_{s}$ are aligned to that of $\mathcal{A}$ through supervision provided by $y_{i}^{s}$ in the source domain and the text supervision provided by $\mathcal{G}$ in the target domain in the form of $z_{i}^{t}$ and $\overline{y}_{i}^{t}$. The source and target datasets are merged ($\hat{\mathcal{X}}=\mathcal{X}_{t}+\mathcal{X}_{s}$) and used for training with a total of $M=N_{s}+N_{t}$ samples. Source samples comprise $(x_{i}^{s},l_{i}^{s},y_{i}^{s})$ triplet, whereas target samples $(x_{i}^{t},l_{i}^{t},\overline{y}_{i}^{t})$ use $\overline{y}_{i}^{t}$ in absence of $y_{i}^{t}$. Target and source images are processed by $\mathcal{V}$ to obtain latent representations $g_{i}^{t}$ and $g_{i}^{s}$, while target captions $l_{i}^{t}$ are processed by $\mathcal{G}$ to obtain latent representations $z_{i}^{t}$ as follows:
@@ 3.4.1 Structure Learning
To ensure the learned visual representations adhere to the structure defined by $\mathcal{A}$, LAGUNA employs a supervised learning approach based on relative encodings. For target samples, we first compute $\mathbf{r}^{z_{i}^{t}}=rel(z_{i}^{t},\mathcal{A})$. For source samples, we compute the relative encoding of the anchor associated with the ground truth class $\mathcal{A}[y_{i}^{s}]$: $\mathbf{r}^{y_{i}^{s}}=rel(\mathcal{A}[y_{i}^{s}],\mathcal{A})$. These encodings represent the relative positions of textual representations with respect to the reference space $\mathcal{A}$ and are not trainable as $\mathcal{G}$ is fixed. On the visual side, we compute the relative encodings of $g_{i}^{t}$ and $g_{i}^{s}$ with respect to the learnable anchors $\mathcal{A}_{t}$ and $\mathcal{A}_{s}$, i.e., $\mathbf{r}^{g_{i}^{t}}=rel(g_{i}^{t},\mathcal{A}_{t})$ and $\mathbf{r}^{g_{i}^{s}}=rel(g_{i}^{s},\mathcal{A}_{s})$. We encourage the visual relative encodings ($\mathbf{r}^{g_{i}^{t}},\mathbf{r}^{g_{i}^{s}}$) to match the text relative encodings ($\mathbf{r}^{z_{i}^{t}},\mathbf{r}^{y_{i}^{s}}$) with an L1 loss:
@@
This loss encourages $g_{i}^{t}$ and $g_{i}^{s}$ to maintain relationships with their respective domain-specific visual anchors that mirror those expressed by $\mathcal{A}$. This process supervises both the learnable anchors $\mathcal{A}_{s},\mathcal{A}_{t}$ and the visual encoder $\mathcal{V}$. However, since this loss alone might lead to the collapse of anchor representations into a small region, hindering classification decision boundaries, we introduce a volume (spread) regularization loss. This loss treats each set of anchors as a multidimensional parallelotope whose volume in the latent space is measured by the determinant of its Gram matrix. Given the Gram matrices for the three sets of anchors:
@@
we encourage the volume of each domain-specific parallelotope to be approximately equal to that of $\mathcal{A}$ with the following loss:
$\mathcal{L}_{Reg}=|logDet(\gamma_{t})-logDet(\gamma)|+|logDet(\gamma_{s})-logDet(\gamma)|.$ (7)
We use log-determinants for numerical stability. This regularization loss ensures that visual anchors occupy a volume similar to that of $\mathcal{A}$, preventing collapse in the representation space.
@@ 3.4.2 Classification Training
We perform classification training concurrently with structure learning. To ensure that source and target visual representations are grounded in the structure imposed by $\mathcal{A}$, while still being free to capture domain-specific nuances, we propose a novel cross-domain attention layer that aims to ground $g_{i}^{t}$ and $g_{i}^{s}$ to the structure of source anchors $\mathcal{A}_{s}$. We include two Cross-Domain attention layers sharing the same weights for the target and source branches of LAGUNA (See Fig. 2). In the target branch, we use $g_{i}^{t}$ as queries, target anchors $\mathcal{A}_{t}$ as keys, and source anchors $\mathcal{A}_{s}$ as values. The output is summed to $g_{i}^{t}$ with a residual connection to include both domain-specific and cross-domain information in the final representation:
@@
Crucially, the cross-domain attention layer constructs an attention map by leveraging the domain-specific relations between the encoded representations and the anchors (i.e., $QK^{T}$). These relations are expected to be similar across domains due to the influence of the structural loss $\mathcal{L}_{S}$. This process results in a unified representation since the output of the attention layer always depends on values coming from the source anchors $\mathcal{A}_{s}$. The residual connection included in the attention layer is introduced to account for domain-dependent class characteristics, which results in the feature vectors $f_{i}^{s}$ and $f_{i}^{t}$. Finally, an MLP classification head processes $f_{i}^{t}$ and $f_{i}^{s}$ to output predictions $\hat{y}_{i}^{t}$ and $\hat{y}_{i}^{s}$. The MLP has shared weights across domains and is optimized with a cross-entropy loss using $y_{i}^{s}$ for source examples and $\overline{y}_{i}^{t}$ in the case of target samples. The overall training objective is:
$\mathcal{L}=\lambda_{1}\mathcal{L}_{CE}(\hat{y_{i}},y_{i}^{*})+\lambda_{2}\mathcal{L}_{S}(\mathbf{r}^{*},\mathbf{r}^{g_{i}})+\lambda_{3}\mathcal{L}_{Reg}(\gamma,\gamma_{*}),$ (10)
@@ 4 Experiments
4.1 Experimental Setup
Datasets. We evaluate LAGUNA on four comprehensive UDA benchmarks: DomainNet (Peng et al., 2018), GeoPlaces (Kalluri et al., 2023) GeoImnet (Kalluri et al., 2023), and Ego2Exo (Kalluri et al., 2024). DomainNet consists of $400K$ images across $345$ classes and is used to evaluate performance in $12$ adaptation settings. GeoImnet and GeoPlaces, subsets of the GeoNet dataset, contain over 750K images, and focus on geographic disparities for objects (GeoImNet with $600$ classes) and scenes (GeoPlaces with $205$ classes). Finally, Ego2Exo is a subset of EgoExo-4D (Grauman et al., 2023) curated for domain adaptation. It involves a total of $9086$ videos and $24$ categories. DomainNet and GeoNet include per-image captions derived from BLIP-2 (Li et al., 2023) and web metadata, respectively, while Ego2Exo provides video descriptions from the original EgoExo-4D.
Model and training details. We use a pre-trained SentenceTransformer (Reimers and Gurevych, 2019) for Stage 1 and a BERT base model (Sanh et al., 2019) as text supervisor (Stage 2). The latter is trained on the source domain of each scenario for 5 epochs with a batch size of 64, a lr of 1e-4, and the AdamW (Loshchilov, 2017) optimizer. The visual encoders (Stage 3) are tailored to each dataset for fair comparisons with existing approaches. For DomainNet, we adapt the Swin-B and ViT-B backbones (Liu et al., 2021b; Dosovitskiy et al., 2020). For GeoImnet and GeoPLaces, we utilize the ViT-B backbone. For Ego2Exo, we employ pre-extracted Omnivore-L features (Grauman et al., 2023). All vision models undergo joint training with the other network components for $10$ epochs with a batch size of $32$, an initial learning rate of 1e41e-4, and cosine scheduling. We set the loss weights to $\lambda_{1}$=1.0, $\lambda_{2}$=0.1, and $\lambda_{3}$=0.001 to normalize the magnitudes of loss components.
@@ 4.1 Experimental Setup
MDD (Zhang et al., 2019) 52.80 41.20 47.80 52.50 42.10 40.70 54.20 54.30 43.10 51.20 43.70 41.70 47.11
SCDA (Li et al., 2021) 54.00 42.50 51.90 55.00 44.10 39.30 53.20 55.60 44.70 56.20 44.10 42.00 48.55
SSRT-B (Sun et al., 2022) 69.90 58.90 66.00 75.80 59.80 60.80 73.20 70.60 62.20 71.40 61.70 55.20 65.41
MemSAC (Kalluri et al., 2022) 63.49 42.14 60.32 72.33 54.92 46.14 73.46 68.04 52.75 74.42 57.79 43.57 59.11
CDTrans (Xu et al., 2021) 66.20 52.90 61.50 72.60 58.10 57.10 72.50 69.00 59.00 72.10 62.90 53.90 63.16
PMTrans (Zhu et al., 2023) 74.10 61.10 70.00 79.30 63.70 62.70 77.50 73.80 62.60 79.80 69.70 61.20 69.63
@@
TextMatch (Kalluri et al., 2024) 71.36 64.30 65.32 81.25 65.65 64.85 81.09 72.65 63.94 81.08 70.84 64.17 70.64
nGramMatch (Kalluri et al., 2024) 68.92 59.82 63.15 76.35 61.72 62.87 76.35 69.28 62.51 76.04 68.52 60.52 67.17
LaGTran (Kalluri et al., 2024) 77.30 68.25 67.35 81.31 67.03 66.81 80.78 75.62 68.08 79.23 73.80 63.44 72.41
@@
ViT-B CLIP* (Radford et al., 2021) 72.39 60.90 66.81 81.37 60.90 66.81 81.37 72.39 66.81 81.37 72.39 69.90 70.38
DAPrompt* (Ge et al., 2025) 73.90 65.90 70.40 84.90 65.80 70.20 84.60 73.50 69.90 84.90 73.80 65.80 73.80
PADCLIP (Lai et al., 2023) 76.40 67.50 72.70 84.20 68.10 71.10 83.60 76.30 71.70 83.50 75.40 67.20 74.81
@@
Table 1: DomainNet dataset results for domain adaptation. Includes 4 domains: Real, Clipart, Sketch, and Painting, for a total of 12 domain adaptation scenarios. The best results are reported in bold while the second best are underlined.
4.2 Results
We benchmark our model against various language-free UDA methods that have reported results on the considered datasets (Zhu et al., 2023; Wei et al., 2021; Kalluri et al., 2022; Zhang et al., 2019; Chen et al., 2022; Xu et al., 2021). Additionally, we also compare with previous approaches using language guidance (Kalluri et al., 2024; Radford et al., 2021; Ge et al., 2025; Lai et al., 2023; Li et al., 2024). Following the evaluation protocols of each benchmark, we report accuracy for DomainNet, GeoImNet, and GeoPlaces, and per-class mean accuracy for Ego2Exo.
DomainNet. Table 1 presents the results on DomainNet, encompassing $12$ UDA scenarios across four distinct domains: Real (R), Clipart (C), Sketch (S), and Painting (P). LAGUNA is compared against previous SOTA methods utilizing the Swin-B and ViT-B backbones as originally adopted by the compared methods to ensure fair comparison. LAGUNA significantly outperforms previous methods in 12 scenarios, with an average performance increase of $3.32\%$ for Swin-B. For ViT-B, LAGUNA leads in 11 of 12 scenarios, boosting performance by an average of $2.51\%$. Importantly, LAGUNA achieves SOTA ViT-B performance without relying on extensive CLIP pretraining and domain-aware prompting, unlike previous models (Lai et al., 2023; Li et al., 2024; Ge et al., 2025).
[…]